# A Closer Look at AUROC and AUPRC under Class Imbalance

### v3

@@ Abstract
In machine learning (ML), a widespread claim is that the area under the precision-recall curve (AUPRC) is a superior metric for model comparison to the area under the receiver operating characteristic (AUROC) for tasks with class imbalance. This paper refutes this notion on two fronts. First, we theoretically characterize the behavior of AUROC and AUPRC in the presence of model mistakes, establishing clearly that AUPRC is not generally superior in cases of class imbalance. We further show that AUPRC can be a harmful metric as it can unduly favor model improvements in subpopulations with more frequent positive labels, heightening algorithmic disparities. Next, we empirically support our theory using experiments on both semi-synthetic and real-world fairness datasets. Prompted by these insights, we conduct a review of over 1.5 million scientific papers to understand the origin of this invalid claim, finding that it is often made without citation, misattributed to papers that do not argue this point, and aggressively over-generalized from source arguments. Our findings represent a dual contribution: a significant technical advancement in understanding the relationship between AUROC and AUPRC and a stark warning about unchecked assumptions in the ML community.
Machine Learning, evaluation, auroc, auprc, Fairness
@@ 1 Introduction
Machine learning (ML), especially in critical domains like healthcare, necessitates carefully selecting and applying evaluation metrics to guide appropriate model choices and understand performance nuances (Hicks et al., 2022). Model evaluation can happen in one of two settings: (1) a methodological/model comparison setting, which occurs outside of a specific deployment setting and target model usage workflows, optimal decision thresholds, or specific false-positive (FP) and false-negative (FN) costs are typically not known, or (2) an application/deployment setting, where reasonably specific estimates of model usage workflows and FP/FN costs can be made. In both of these settings, appropriate evaluation metric choice is critical—inappropriate evaluation metrics can hinder innovation when used for model comparison and can lead to significant real world costs (e.g., mis-diagnosis in a medical setting) in deployment settings.
This study focuses on two pivotal metrics for binary classification tasks that are widely used in both evaluation contexts: the Area Under the Precision-Recall Curve (AUPRC) and the Area Under the Receiver Operating Characteristic (AUROC). Central to this paper is the following key claim:
@@
Claim 1 is made widely in both the scientific literature (Wagner et al., 2023; Choi et al., 2018; Hsu et al., 2020; Gong et al., 2021) and in popular press sources (Czakon, 2022; Mazzanti, 2023) and has been justified on numerous grounds (See sources collated in the literature review in Section 5). Despite this, we show in this work that this claim is, in fact, wrong, and many of its justifications are invalid or misapplied in common ML settings. More specifically, we show the following:
1) AUROC and AUPRC only differ with respect to model-dependent parameters in that AUROC weighs all false positives equally, whereas AUPRC weighs false positives at a threshold $\tau$ with the inverse of the model’s likelihood of outputting any scores greater than $\tau$ (Theorem 1). This result shows that we can reason about the suitability of optimizing or selecting by AUROC vs. AUPRC on the basis of whether we care more about reducing false positives above low thresholds or high thresholds. In particular,
@@
3) AUPRC can unduely prioritize improvements to higher-prevalence subpopulations at the expense of lower-prevalence subpopulations, raising serious fairness concerns in any multi-population use cases (Theorem 3).
@@
In this work, we will establish these three claims theoretically, via synthetic experiments, and with real-world validation on popular public fairness datasets. We will additionally demonstrate through an extensive, large-language model aided literature review of over 1.5 million scientific papers, that Claim 1 has been used to motivate numerous improper uses of AUPRC relative to AUROC across high-impact domains like healthcare in a range of well known venues, including Cancer Cell, Nature Scientific Reports, AAAI, and NeurIPS. Through this paper, we hope to shed greater light on the nuances of appropriate evaluation and provide key guidance to limit future misuse of evaluation metrics in the scientific and machine learning communities.
@@
Figure 1: a) Consider a model $f$ yielding continuous output scores for a binary classification task applied to a dataset consisting of two distinct subpopulations, $\mathcal{A}\in\{0,1\}$ (e.g., race). If we order samples in ascending order of output score, we define a mistake to be when two neighboring samples are out-of-order with respect to the classification label (such as Mistake 1, 2, 3, and 4 in the figure). Each such mistake represents an unambiguous, indivisible opportunity for model improvement; namely, swap the out-of-order samples, and the model will improve. Under any such swapping operation, the model’s AUROC will improve by the same amount, while the model’s AUPRC will improve by an amount correlated with the inverse of the number of samples the model assigns a higher score than the samples in the mistake (Theorem 2). b) If we are performing model comparison outside of a deployment setting, the we will have no insight into the expected costs of various kinds of mistakes and we have no reason to believe we will primarily be concerned with samples $x$ that are drawn from any particular region in the model output space. Therefore, in this setting, an appropriate model evaluation metric should improve by the same amount regardless of which mistake is corrected. c) In a deployment setting where false negatives have a high cost relative to false positives, such as a routine screening test for a serious disease like cancer, then an appropriate evaluation metric should optimize for ensuring that we can most efficiently catch all cases of the disease and minimize the number of false negatives. This corresponds to favoring mistakes that have lower scores, regardless of the fact that many such tasks have significant rates of class imbalance. d) In a deployment setting where a limited number of key resources will be distributed among members of a large population, in a manner that requires certain subpopulations to be offered commensurate possible benefit from the intervention for ethical reasons, a proper evaluation metric should prioritize the importance of within-group, high-score mistakes such that the highest risk members of each subgroup can receive the limited number of interventions. e) In a deployment setting where false positives are expensive relative to false negatives and there are no fairness concerns, such as prioritizing small molecule leads for drug development, where the top $k$ most likely molecules will be presented for expensive follow-up experimental validation, appropriate evaluation should favor model improvements in decreasing order with score, to maximize the likelihood of successfully identifying appropriate molecules among the limited set of $k$ selections.
@@ 2 Theoretical Analyses
All notation used is defined in Appendix Section B.
@@ 2.1 Relationship Between AUROC and AUPRC
Let $\mathcal{X},\mathcal{Y}={0,1}$ represent a paired feature and binary classification label space from which i.i.d. samples $(x,y)\in\mathcal{X}\times\mathcal{Y}$ are drawn via the joint distribution over the random variables $\mathsf{x},\mathsf{y}$. Let $f:\mathcal{X}\to(0,1)$ be a binary classification model outputting continuous probability scores over this space. Then,
@@
$\mathrm{AUPRC}(f)$ $=1-p_{\mathsf{y}}(0)\mathbb{E}_{t\sim f(\mathsf{x})|\mathsf{y}=1}\left[\frac{\mathrm{FPR}(f,t)}{P(f(\mathsf{x})>t)}\right]$
We provide the proof in Appendix Section C. The two key intuitions are that integrating over the TPR is equivalent to taking the expectation over the induced distribution of positive sample scores, and that via Bayes rule, $\mathrm{Prec}(f,\tau)=1-p_{\mathsf{y}}(0)\frac{\mathrm{FPR}(f,\tau)}{P(f(\mathsf{x})>\tau)}$.
Despite its simplicity, Theorem 1 has far-reaching implications. Namely, it reveals that the only difference between AUROC and AUPRC with respect to model dependent parameters (i.e., omitting the dependence of AUPRC on the fixed prevalence of the dataset, which is not model varying) is that optimizing AUROC equates to minimizing the expected false positive rate over all positive samples in an unweighted manner (equivalently, in expectation over the distribution of positive sample scores) whereas optimizing AUPRC equates to minimizing the expected false positive rate over all positive samples weighted by the inverse of the model’s “firing rate” ($P(f(\mathsf{x})>\tau)$) at the given positive sample score. This preference can be crystallized when we examine how AUROC vs. AUPRC would prioritize correcting indivisible units of model improvements, termed “mistakes” which we will discuss next.
@@ 2.2 AUPRC prioritizes high-score mistakes, AUROC treats all mistakes equally
Understanding how a given evaluation metric prioritizes the correction of various kinds of model mistakes or errors offers significant insight into when that metric should be used for optimization or model selection. To examine this topic for AUROC and AUPRC, consider the following definition of an “incorrectly ranked adjacent pair”, which we will colloquially refer to as a “model mistake”:
@@
Let $f,\mathcal{X},\mathcal{Y},\mathsf{x},\mathsf{y}$ be defined as in Theorem 1. Further, let us suppose we have sampled a static dataset from $\mathsf{x},\mathsf{y}$ for evaluation which will be denoted $\boldsymbol{X},\boldsymbol{y}=\{(x_{1},y_{1}),\ldots,(x_{N},y_{N})\}$, for $x_{i}\in\mathcal{X},y_{i}\in\{0,1\},$ and $N\in\mathbb{N}$. We assume for convenience that $f$ is an injective map and all $x_{i}$ are distinct (i.e., $\forall(i,j):x_{i}\neq x_{j}$ which, by injectivity of $f$, implies that $f(x_{i})\neq f(x_{j})$).
@@
$f_{i}^{\prime}:\begin{cases}f(x)&\text{if }x\notin\{x_{i},x_{i+1}\}\\ f(x_{i+1})&\text{if }x=x_{i}\\ f(x_{i})&\text{if }x=x_{i+1}.\\ \end{cases}$
@@
The proof for Theorem 2 can be found in Appendix D. This proof simply stems from the fact that correcting a single mistake $(x_{i},x_{j})$ (as defined in Definition 2.1) always changes the false positive rate by the same amount, and only changes it at the threshold $f(x_{i})$. This, combined with the formalization of AUROC and AUPRC in Theorem 1, establishes the proof. Note that this Theorem can be trivially extended to include a case where ties are possible simply by noting that “swapping” two samples $x_{i}$ and $x_{j}$ in the manner of the theorem results in no change to either AUROC or AUPRC, and similarly by the same reasoning separating any tie in the appropriate direction will improve AUROC uniformly over samples and will improve AUPRC in a manner monotonic with model score.
@@ 2.3 AUPRC is explicitly discriminatory in favor of high-scoring subpopulations
Let $f,\mathcal{X},\boldsymbol{X},\boldsymbol{y},N,M,$ and $f^{\prime}_{j}$ all be defined as in Theorem 2. Further, suppose that in this setting the domain $\mathcal{X}$ now contains an attribute defining two subgroups, $\mathcal{A}=\{0,1\}$, such that for any sample $(x_{i},y_{i})$, $a_{i}$ denotes the subgroup to which that sample belongs. Let $f$ be perfectly calibrated for samples in subgroup $a=0$, such that $P(y=1|a=0,f(x)=t)=t$. Let $p_{0}=p(y=1|a=0)$ denote the prevalence of the label $y$ over subgroup $a=0$. Then,
$\lim_{p_{0}\to 0}P\left(a_{i}=a_{i+1}=1\middle|i=\operatorname*{arg\,max}_{j\in M}\left(\mathrm{AUPRC}(f_{j}^{\prime})\right)\right)=1.$
Essentially, Theorem 3 (proof provided in Appendix E) shows that for any model of interest, provided the model is calibrated, there exists a prevalence disparity sufficiently severe such that the likelihood of the mistake occurring with highest score (which will maximally improve AUPRC) belonging to anything other than the high prevalence subgroup goes to zero. This demonstrates that AUPRC provably favors higher prevalence subpopulations under sufficiently severe class imbalance.
Note that this property is, generally speaking, not desirable. In particular, this property establishes that in settings where model fairness among a set of subpopulations in the data is important, AUPRC should not be used as an evaluation metric due to the risk that it will introduce biases in favor of the highest prevalence subpopulations. We validate this result empirically over both synthetic and real-world data in Section 3, demonstrating that the import of Theorem 3 is not merely limited to an analytical curiosity but can have real world impact on algorithmic disparities in practice.
Furthermore, note that this theorem does not indicate that AUPRC will be superior to AUROC for differentiating a low prevalence (or low risk) subpopulation relative to a high-risk subpopulation, a property that is sometimes attributed to AUPRC in the literature. Rather, Theorem 3 shows that maximizing AUPRC will be more likely to optimize solely within the high-risk subgroup, rather than optimizing to differentiate across subgroups, as low-risk subgroup samples will only occur in lower-score regions under severe class imbalance.
@@ 3.1 Synthetic Optimization Experiments Demonstrate AUPRC-induced Disparities
Let $Y\in\{0,1\}$ be the binary label, $S\in[0,1]$ be the predicted score, and $G\in\{G_{1},G_{2}\}$ be the subpopulation. We fix $\mathbb{P}(Y=1|G=G_{1})=0.05$ and $\mathbb{P}(Y=1|G=G_{2})=0.01$. We sample a dataset for each group $\mathcal{D}_{g}=\{(S_{1},Y_{1}),...,(S_{n_{g}},Y_{n_{g}})\}$, such that $\text{AUROC}(\mathcal{D}_{1})\approx\text{AUROC}(\mathcal{D}_{2})\approx\text{AUROC}(\mathcal{D}_{1}\cup\mathcal{D}_{2})=0.85$ (A target AUROC of 0.65 was also profiled in Appendix Figure 5).
@@
Our main experimental challenge is to determine how to simulate “optimizing” or “selecting” a model by AUROC or AUPRC. We explore two approaches here. First, we can simply correct the atomic mistake that maximally improves AUROC or AUPRC in each optimization iteration. In our experiments, we use $n_{1}=n_{2}=200$ and optimize for $50$ steps for this experiment. This is the most straightforward optimization procedure to analyze, but it is unrealistic. In real optimization scenarios, larger model changes will be made at once, and a model will have an opportunity to degrade performance in some regions in order to improve it in others.
@@
Next, we profile an optimization procedure that randomly permutes all the (sorted) model scores up to 3 positions. This has the effect of randomly adjusting all model scores, and can worsen model performance under some random permutations, but offset precisely the same capacity to the low and high prevalence subgroups. To ensure the model is under some optimization constraint (and therefore does not always find the “perfect” permutation to maximize both metrics identically), we allow the model to sample only 15 possible permutations before choosing the best option. This means the system will be forced to navigate optimization trade-offs between which permutations improve the right regions of the score most effectively among its limited set. We use $n_{1}=n_{2}=100$ for these experiments and optimize for $25$ total steps.
@@
Figure 2: Synthetic experiment per-group AUROC after successively either fixing individual mistakes, as defined in Definition 2.1, (a) and b)) or successively choosing the optimal score permutation (c) and d)) in order to optimize either AUROC (a) and c)) or AUPRC (b) and d)). It is clear across both forms of optimization that AUPRC definitively favors the higher prevalence subpopulation, whereas AUROC treats subgroups approximately equally. Similar patterns were observed when comparing per-group AUPRCs over the same experimental procedures, as shown in Appendix Figure 4.
@@
Across both settings, we run these experiments across 20 randomly sampled datasets and show the mean and an empirical 90% confidence interval around the mean in Figure 2. We present a formal mathematical formulation of these perturbations, as well as profile a third random perturbation method, in Appendix F.3.

### v4

@@ Abstract
In machine learning (ML), a widespread claim is that the area under the precision-recall curve (AUPRC) is a superior metric for model comparison to the area under the receiver operating characteristic (AUROC) for tasks with class imbalance. This paper refutes this notion on two fronts. First, we theoretically characterize the behavior of AUROC and AUPRC in the presence of model mistakes, establishing clearly that AUPRC is not generally superior in cases of class imbalance. We further show that AUPRC can be a harmful metric as it can unduly favor model improvements in subpopulations with more frequent positive labels, heightening algorithmic disparities. Next, we empirically support our theory using experiments on both semi-synthetic and real-world fairness datasets. Prompted by these insights, we reviewed over 1.5 million scientific papers to understand the origin of this invalid claim–finding it is often made without citation, misattributed to papers that do not argue this point, and aggressively overgeneralized from source arguments. Our findings represent a dual contribution: a significant technical advancement in understanding the relationship between AUROC and AUPRC and a stark warning about unchecked assumptions in the ML community.
@@ 1 Introduction
Machine learning (ML), especially in critical domains like healthcare, necessitates careful selection and application of evaluation metrics to guide appropriate model choices and understand performance nuances [150]. Model evaluation can happen in one of two settings: (1) a methodological/model comparison setting, which occurs outside of a specific deployment setting and in which target model usage workflows, optimal decision thresholds, or specific false-positive (FP) and false-negative (FN) costs are typically not known, or (2) an application/deployment setting, where reasonably specific estimates of model usage workflows and FP/FN costs can be made. In both settings, appropriate metric choice is critical, as inappropriate selection can hinder innovation when used for model comparison and lead to significant real-world costs (e.g., misdiagnosis in a medical setting) in deployment settings.
This study focuses on two widely used metrics for binary classification tasks across both evaluation contexts: Area Under the Precision-Recall Curve (AUPRC) and Area Under the Receiver Operating Characteristic (AUROC). Central to this paper is the following key claim:
@@
Claim 1 is made widely in both the scientific literature [399, 71, 159, 124], in ML educational content [119, 141], and in popular press sources [80, 254]. It is so widespread that even basic search results for queries relating to AUROC and AUPRC [^1] See https://archive.is/qXPKu, which shows results dominated by those making Claim 1 and large language model assistants like ChatGPT or Github Co-pilot will profess its veracity. [^2] See https://chat.openai.com/share/f8f7fddb-1553-41a5-976d-789e2f3a90d6 Throughout these sources, it has been justified on numerous, often imprecise grounds (see Section 5), but despite this extensive attention, we show in this work that this claim is, in fact, wrong, and may be dangerous from a model fairness perspective; further, many of its justifications are invalid or misapplied in common ML settings. More specifically, we show the following:
1) AUROC and AUPRC only differ with respect to model-dependent parameters in that AUROC weighs all false positives equally, whereas AUPRC weighs false positives at a threshold $\tau$ with the inverse of the model’s likelihood of outputting any scores greater than $\tau$ (Theorem 1). This result shows that we can reason about the suitability of AUROC vs. AUPRC based on whether we care more about reducing false positives above low thresholds or high thresholds. In particular,
@@
3) AUPRC can unduly prioritize improvements to higher-prevalence subpopulations at the expense of lower-prevalence subpopulations, raising serious fairness concerns in any multi-population use cases (Theorem 3).
@@
In this work, we establish these three claims both theoretically and empirically via synthetic experiments and real-world validation on popular public fairness datasets. In addition, we demonstrate through an extensive, large-language model aided literature review of over 1.5 million scientific papers that Claim 1 has been used to motivate numerous improper uses of AUPRC relative to AUROC across high-stakes domains like healthcare and in several established venues, including AAAI, NeurIPS, ICML, ICLR, Cancer Cell, Nature Journals, PNAS, and more. Through this paper, we hope to shed light on the nuances of appropriate evaluation and provide key guidance to limit future misuse of evaluation metrics in the scientific and machine learning communities.
@@
Figure 1: a) Consider a model $f$ yielding continuous output scores for a binary classification task applied to a dataset consisting of two distinct subpopulations, $\mathcal{A}\in\{0,1\}$. If we order samples in ascending order of output score, each misordered pair of samples (e.g., mistake 1-4) represents an opportunity for model improvement. Theorem 2 shows that a model’s AUROC will improve by the same amount no matter which mistake you fix, while the model’s AUPRC will improve by an amount correlated with the score of the sample. b) When comparing models absent a specific deployment scenario, we have no reason to value improving one mistake over another, and model evaluation metrics should therefore improve equally regardless of which mistake is corrected. c) When false negatives have a high cost relative to false positives, evaluation metrics should favor mistakes that have lower scores, regardless of any class imbalance. d) When limited resources will be distributed among a population according to model score, in a manner that requires certain subpopulations to all be offered commensurate possible benefit from the intervention for ethical reasons, evaluation metrics should prioritize the importance of within-group, high-score mistakes such that the highest risk members of all subgroups receive interventions. e) When false positives are expensive relative to false negatives and there are no fairness concerns, evaluation metrics should favor model improvements in decreasing order with score.
@@ 2 Theoretical Analyses
Please note that all notation used is defined in Appendix Section C.
@@ 2.1 Relationship between AUROC and AUPRC
Let $\mathcal{X},\mathcal{Y}=\{0,1\}$ represent a paired feature and binary classification label space from which i.i.d. samples $(x,y)\in\mathcal{X}\times\mathcal{Y}$ are drawn via the joint distribution over the random variables $\mathsf{x},\mathsf{y}$. Let $f:\mathcal{X}\to(0,1)$ be a binary classification model outputting continuous probability scores over this space. Then,
@@
$\mathrm{AUPRC}(f)$ $=1-P_{\mathsf{y}}(y=0)\mathbb{E}_{t\sim f(\mathsf{x})|\mathsf{y}=1}\left[\frac{\mathrm{FPR}(f,t)}{P_{\mathsf{x}}(f(x)>t)}\right]$
We provide the proof in Appendix Section D. The two key intuitions are that integrating over the TPR is equivalent to taking the expectation over the induced distribution of positive sample scores, and that via Bayes rule, $\mathrm{Prec}(f,\tau)=1-P_{\mathsf{y}}(y=0)\frac{\mathrm{FPR}(f,\tau)}{P_{\mathsf{x}}(f(x)>\tau)}$.
Despite its simplicity, Theorem 1 has far-reaching implications. Namely, it reveals that the only difference between AUROC and AUPRC with respect to model dependent parameters (i.e., omitting the dependence of AUPRC on the fixed prevalence of the dataset, which is not model varying) is that optimizing AUROC equates to minimizing the expected false positive rate over all positive samples in an unweighted manner (equivalently, in expectation over the distribution of positive sample scores) whereas optimizing AUPRC equates to minimizing the expected false positive rate over all positive samples weighted by the inverse of the model’s “firing rate” ($P_{\mathsf{x}}(f(x)>\tau)$) at the given positive sample score. This preference can be crystallized when we examine how AUROC vs. AUPRC would prioritize correcting indivisible units of model improvements, termed “mistakes” which we will discuss next.
@@ 2.2 AUPRC prioritizes high-score mistakes, AUROC treats all mistakes equally
Understanding how a given evaluation metric prioritizes correcting various kinds of model mistakes or errors offers significant insight into when that metric should be used for optimization or model selection. To examine this topic for AUROC and AUPRC, consider the following definition of an “incorrectly ranked adjacent pair”, which we will colloquially refer to as a “model mistake”:
@@
Let $f,\mathcal{X},\mathcal{Y},\mathsf{x},\mathsf{y}$ be defined as in Theorem 1. Further, let us suppose we have sampled a static dataset from $\mathsf{x},\mathsf{y}$ for evaluation which will be denoted $\boldsymbol{X},\boldsymbol{y}=\{(x_{1},y_{1}),\ldots,(x_{N},y_{N})\}$, for $x_{i}\in\mathcal{X},y_{i}\in\{0,1\},$ and $N\in\mathbb{N}$. We assume for convenience that $f$ is an injective map and all $x_{i}$ are distinct (i.e., $\forall(i,j)|i\neq j:x_{i}\neq x_{j}$ which, by injectivity of $f$, implies that $f(x_{i})\neq f(x_{j})$).
@@
$f_{i}^{\prime}:x\mapsto\begin{cases}f(x)&\text{if }x\notin\{x_{i},x_{i+1}\}\\ f(x_{i+1})&\text{if }x=x_{i}\\ f(x_{i})&\text{if }x=x_{i+1}.\\ \end{cases}$
@@
The proof for Theorem 2 can be found in Appendix E. This proof simply stems from the fact that correcting a single mistake $(x_{i},x_{j})$ (as defined in Definition 2.1) always changes the false positive rate by the same amount, and only changes it at the threshold $f(x_{i})$. This, combined with the formalization of AUROC and AUPRC in Theorem 1, establishes the proof. Note that this Theorem can be trivially extended to include a case where ties are possible simply by noting that “swapping” two samples $x_{i}$ and $x_{j}$ where $f(x_{i})=f(x_{j})$ in the manner of the theorem results in no change to either AUROC or AUPRC, and similarly by the same reasoning separating any tie in the appropriate direction will improve AUROC uniformly over samples and will improve AUPRC in a manner monotonic with model score.
@@ 2.3 AUPRC is explicitly discriminatory in favor of high-scoring subpopulations
Let $f,\mathcal{X},\boldsymbol{X},\boldsymbol{y},N,M,$ and $f^{\prime}_{j}$ all be defined as in Theorem 2. Further, suppose that in this setting the domain $\mathcal{X}$ now contains an attribute defining two subgroups, $\mathcal{A}=\{0,1\}$, such that for any sample $(x_{i},y_{i})$, $a_{i}$ denotes the subgroup to which that sample belongs. Let $f$ be perfectly calibrated for samples in subgroup $a=0$, such that $P_{\mathsf{y}|\mathsf{a},\mathsf{x}}(y=1|a=0,f(x)=t)=t$. Then,
$\lim_{P_{\mathsf{y}|\mathsf{a}}(y=1|a=0)\to 0}P\left(a_{i}=a_{i+1}=1\middle|i=\operatorname*{arg\,max}_{j\in M}\left(\mathrm{AUPRC}(f_{j}^{\prime})\right)\right)=1.$
Essentially, Theorem 3 (proof provided in Appendix F) shows the following. Suppose we are training a model $f$ over a dataset with two subpopulations: Population $a=0$ and $a=1$. If the model $f$ is calibrated and the rate at which $y=1$ for population $a=0$ is sufficiently low relative to the rate at which $y=1$ for population $a=1$, then the mistake that, were it fixed, would maximally improve the AUPRC of $f$ will be a mistake purely in population $a=1$. This demonstrates that AUPRC provably favors higher prevalence subpopulations (those with a higher base rate at which $y=1$) under sufficiently severe prevalence imbalance between subpopulations.
Note that this property is, generally speaking, not desirable. In particular, this property establishes that in settings where model fairness among a set of subpopulations in the data is important, AUPRC should not be used as an evaluation metric due to the risk that it will introduce biases in favor of the highest prevalence subpopulations. We validate this result empirically over both synthetic and real-world data in Section 3, demonstrating that the import of Theorem 3 is not merely limited to an analytical curiosity but can have real-world impact on algorithmic disparities in practice. Furthermore, note that this theorem does not indicate that AUPRC will be superior to AUROC for differentiating a low prevalence (or low risk) subpopulation relative to a high-risk subpopulation, a property that is sometimes attributed to AUPRC in the literature. Rather, Theorem 3 shows that maximizing AUPRC will be more likely to optimize solely within the high-risk subgroup, rather than optimizing to differentiate across subgroups, as low-risk subgroup samples will predominantly occur in lower-score regions under severe class imbalance.
@@ 3.1 Synthetic optimization experiments demonstrate AUPRC-induced disparities
In this section, we use a carefully constructed synthetic optimization procedure to demonstrate that, when all other factors are equal, optimizing by or performing model selection on the basis of AUPRC vs. AUROC risks excacerbating algorithmic disparities in the manner predicted by Theorem 3. For analyses under more realistic conditions with more standard models, see our real-world experiments in Section 3.2.
@@
Let $y\in\{0,1\}$ be the binary label, $s\in[0,1]$ be the predicted score, and $a\in\{1,2\}$ be the subpopulation. We fix $P_{\mathsf{y}|\mathsf{a}}(y=1|a=1)=0.05$ and $P_{\mathsf{y}|\mathsf{a}}(y=1|a=2)=0.01$. We sample a dataset for each group $\mathcal{D}_{a}=\{(s_{1},y_{1}),...,(s_{n_{a}},y_{n_{a}})\}$, such that $\text{AUROC}(\mathcal{D}_{1})\approx\text{AUROC}(\mathcal{D}_{2})\approx\text{AUROC}(\mathcal{D}_{1}\cup\mathcal{D}_{2})=0.85$ (See Appendix G.1; A target AUROC of 0.65 was also profiled in Appendix Figure 5).
@@
Our main experimental challenge is to determine how to simulate “optimizing” or “selecting” a model by AUROC or AUPRC. Simulating optimizing by these metrics allows us to explicitly assess how the use of either AUPRC or AUROC as an evaluation metric in model selection processes such as hyperparameter tuning or architecture search, can translate into model-induced inequities in dangerous ways. We explore two approaches here. First, we can simply correct the atomic mistake that maximally improves AUROC or AUPRC in each optimization iteration. In our experiments, we use $n_{1}=n_{2}=200$ and optimize for $50$ steps for this experiment. This is the most straightforward optimization procedure to analyze, but it is unrealistic. In real optimization scenarios, larger model changes will be made at once, and a model will have an opportunity to degrade performance in some regions in order to improve it in others.
@@
Next, we profile an optimization procedure that randomly permutes all the (sorted) model scores up to 3 positions (See Appendix G.3 for details). This has the effect of randomly adjusting all model scores, and can worsen model performance under some random permutations, but offers precisely the same “optimization capacity” to the low and high prevalence subgroups. To ensure the model is under some optimization constraint (and therefore does not always find the “perfect” permutation to maximize both metrics identically), we allow the model to sample only 15 possible permutations before choosing the best option. This means the system will be forced to navigate optimization trade-offs between which permutations improve the right regions of the score most effectively among its limited set. We use $n_{1}=n_{2}=100$ for these experiments and optimize for $25$ total steps.
@@
Figure 2: Synthetic experiment per-group AUROC, showing a confidence interval spanning the 5th to 95th percentile of results observed across all seeds, after successively either fixing individual mistakes, as defined in Definition 2.1, (a) and b)) or successively choosing the optimal score permutation (c) and d)) in order to optimize either AUROC (a) and c)) or AUPRC (b) and d)). It is clear across both forms of optimization that AUPRC definitively favors the higher prevalence subpopulation, whereas AUROC treats subgroups approximately equally. Similar patterns were observed when comparing per-group AUPRCs over the same experimental procedures, as shown in Appendix Figure 4.
@@
Across both settings, we run these experiments across 20 randomly sampled datasets and show the mean and an empirical 90% confidence interval around the mean in Figure 2. We present a formal mathematical formulation of these perturbations, as well as profile a third random perturbation method, in Appendix G.3.